OpenStudy (anonymous):

People please help! If 24 gallons of water are poured into an empty tank , then 3/4 of the tank is filled. How many gallons does the full tank hold?

OpenStudy (klimenkov):

\(x\cdot\frac34=24\) Solve it and get an answer.

OpenStudy (anonymous):

What is the answer?

OpenStudy (anonymous):

Ohh I forgot I need to explain my work too!!

OpenStudy (anonymous):

Help me please this is 25% of my grade

OpenStudy (anonymous):

well since you've been given the equation.. you can say that the quantity of gallons needed to fill the tank is x. Since 3/4 of the tank is full with 24 gallons you get 3/4x = 24

OpenStudy (anonymous):

then the total of the tank is solving for x. x = 4/3 * 24

OpenStudy (anonymous):

x = 32. you plug that into the equation and you get that 3/4 of the tank is full at 24 gallons if the tank can hold a total of 32 gallons.

OpenStudy (anonymous):

let the total capacity of the tank be = x gallons then 3/4 of x is 24 i.e. 3/4 *x = 24 => x = (24*4)/3 = 96/3 = 32 gallons so the tank can hold 32 gallons

OpenStudy (anonymous):

Oh thanks guys !!